# To Judge or not to Judge: Using LLM Judgements for Advertiser Keyphrase Relevance at eBay

### v2

@@ Abstract
E-commerce sellers are recommended keyphrases based on their inventory on which they advertise to increase buyer engagement (clicks/sales). The relevance of advertiser keyphrases plays an important role in preventing the inundation of search systems with numerous irrelevant items that compete for attention in auctions, in addition to maintaining a healthy seller perception. In this work, we describe the shortcomings of training Advertiser keyphrase relevance filter models on click/sales/search relevance signals and the importance of aligning with human judgment, as sellers have the power to adopt or reject said keyphrase recommendations. In this study, we frame Advertiser keyphrase relevance as a complex interaction between 3 dynamical systems — seller judgment, which influences seller adoption of our product, Advertising, which provides the keyphrases to bid on, and Search, who holds the auctions for the same keyphrases. This study discusses the practicalities of using human judgment via a case study at eBay Advertising and demonstrate that using LLM-as-a-judge en-masse as a scalable proxy for seller judgment to train our relevance models achieves a better harmony across the three systems — provided that they are bound by a meticulous evaluation framework grounded in business metrics.
@@ 1 Introduction
In the realm of e-commerce, sellers adeptly employ online advertising solutions, including the use of keyphrase recommendations, to counteract their often suboptimal positions in organic search results. This approach enables them to secure a beneficial presence on the search results page (SRP) and to cultivate heightened engagement with potential buyers. At eBay, the process of Advertiser Keyphrase recommendation comprises three fundamental steps: retrieval, which entails the aggregation of keyphrases that are pertinent to the product; relevance, which involves the further refinement of these keyphrases to ascertain their pertinence to the item in question; and ranking, which prioritizes the adoption of keyphrases based on the seller’s budgetary constraints and preferences. The role of keyphrase relevance is of paramount importance, as it shapes seller perceptions and aids in preventing the inundation of search systems with numerous irrelevant items vying for attention in auctions. Advertiser keyphrase relevance models are trained to discern overarching patterns within click and sales data. A keyphrase that has attracted a significant amount of clicks or sales relative to an item can be deemed relevant to that item. Fundamentally, clicks and sales constitute robust affirmative indicators of relevance; however, they are not effective as indicators of non-relevance. E-commerce datasets are susceptible to missing-not-at-random (MNAR) conditions due to various biases  [3, 34, 16, 39, 17, 5, 28, 20]. The absence of clicks for an item in relation to a specific query does not inherently imply the item’s irrelevance. Within the e-commerce framework, buyers function as annotators, yet, unlike traditional annotators, they are exposed to a biased presentation of items as a result of search rankings, which in turn influences their annotations in terms of clicks or sales. Consequently, an item that lacks popularity will occupy a suboptimal position on the SRP because search rankings are derived from clicks, and as a result, it may fail to secure any clicks or sales. This bias undermines the reliability of negative relevance signals that are based solely on clicks or sales.
@@
The interaction between the Sellers, Advertising and Search displays an inherent imbalance — Sellers have the ultimate say in adopting Advertising’s keyphrases while Search has the ultimate say in rejecting those keyphrases at the time of auction. For instance, take Itm1 from Figure 3; Advertising considers KP1 and KP4 irrelevant, while eBay Search regards KP2, KP3, and KP5 as irrelevant and the seller rejecting KP2 as a legit keyphrase, leading to Itm1 not participating in any auctions. This outcome occurs regardless of the true relevance of the keyphrases, meaning that even if KP2, KP3, and KP5 are genuinely relevant, Itm1 will not enter those auctions because the seller and eBay Search does not find them relevant. An ideal scenario of complete agreement would be for Itm5 and Itm3 where all parties are in complete agreement.  [^1] For Itm3 the outcome is not as great as Itm5 as there are still some keyphrases which are not getting into auctions. However, that is more of a retrieval problem than a relevance problem, i.e., why are the keyphrases KP2and KP4 getting recommended in the first place?
The comprehension of auction dynamics and mechanisms is paramount for the successful execution of advertising campaigns. However, this understanding is further complicated by the complex nature of sellers’ perceptions. Promoted Listings Priority, our product, enables manual targeting, which allows sellers to select their preferred keyphrases while also granting them the option to follow our recommendations. For instance, should sellers judge KP2 for Itm1 as irrelevant, they may choose to disregard this particular keyphrase, thereby rendering our suggestions ineffective. This scenario can escalate, as sellers often encounter an abundance of "irrelevant" keyphrases, discouraging them from proceeding with the adoption of our product. Such outcomes not only lead to diminished Seller Satisfaction scores, but also represent a wasted utilization of resources and opportunities for eBay. Moreover, offering ostensibly "sensible" keyphrases lacking substantial targeting capabilities can result in subpar advertiser campaign performance, potentially prompting sellers to leave the platform. Consequently, it is imperative to ensure that keyphrases resonate with human judgment while simultaneously maintaining robust performance. In this study, we interpret Advertiser keyphrase relevance as a multifaceted interplay among three dynamical systems: seller judgment impacting their product adoption, Advertising supplying the bid and the keyphrases, and Search orchestrating auctions for those keyphrases.
@@ 3.1.2 Click Data
Since clicks are a strong indicator for relevance, we gather click data amounting to roughly 24 million item keyphrase pairs over a period of 30 days with at least 30 impressions and 1 clicks with a minimum CTR of 0.1. Click data can be seen as data annotated by millions of buyers, and it is a stronger signal as buyers are spending their time to do so. The minimum thresholds of CTR, impressions and clicks are used to eliminate noise such as 1 click 1 impression or 1 click and 1000 impression. An important thing to note is that due to the nature of the auction process, all click data is deemed relevant by eBay Search (i.e. they are only surfaced to buyers because it is deemed relevant by eBay Search).

### v3

@@ Abstract
E-commerce sellers are recommended keyphrases based on their inventory on which they advertise to increase buyer engagement (clicks/sales). The relevance of advertiser keyphrases plays an important role in preventing the inundation of search systems with numerous irrelevant items that compete for attention in auctions, in addition to maintaining a healthy seller perception. In this work, we describe the shortcomings of training Advertiser keyphrase relevance filter models on click/sales/search relevance signals and the importance of aligning with human judgment, as sellers have the power to adopt or reject said keyphrase recommendations. In this study, we frame Advertiser keyphrase relevance as a complex interaction between 3 dynamical systems — seller judgment, which influences seller adoption of our product, Advertising, which provides the keyphrases to bid on, and Search, who holds the auctions for the same keyphrases. This study discusses the practicalities of using human judgment via a case study at eBay Advertising and demonstrates that using LLM-as-a-judge en-masse as a scalable proxy for seller judgment to train our relevance models achieves a better harmony across the three systems, provided that they are bound by a meticulous evaluation framework grounded in business metrics.
@@ 1 Introduction
In the realm of e-commerce, sellers adeptly employ online advertising solutions, including the use of keyphrase recommendations, to counteract their often suboptimal positions in organic search results. This approach enables them to secure a beneficial presence on the search results page (SRP) and cultivate a greater engagement with potential buyers, see Figure 1. At eBay, the process of Advertiser keyphrase recommendation comprises three fundamental steps: retrieval, which entails the aggregation of keyphrases that are pertinent to the product; relevance, which involves the further refinement of these keyphrases to ascertain their pertinence to the item in question; and ranking, which prioritizes the adoption of keyphrases based on the seller’s budgetary constraints and preferences. The role of keyphrase relevance is of paramount importance, as it shapes seller perceptions and aids in preventing the inundation of search systems with numerous irrelevant items vying for attention in auctions. Advertiser keyphrase relevance models are trained to discern overarching patterns within click and sales data. A keyphrase that has attracted a significant amount of clicks or sales relative to an item can be deemed relevant to that item. Fundamentally, clicks and sales constitute robust affirmative indicators of relevance; however, they are not effective as indicators of non-relevance. E-commerce datasets are susceptible to missing-not-at-random (MNAR) conditions due to various biases  [3, 35, 16, 40, 17, 5, 29, 20]. The absence of clicks for an item in relation to a specific query does not inherently imply the item’s irrelevance. Within the e-commerce framework, buyers function as annotators, yet, unlike traditional annotators, they are exposed to a biased presentation of items as a result of search rankings, which in turn influences their annotations in terms of clicks or sales. Consequently, an item that lacks popularity will occupy a suboptimal position on the SRP because search rankings are derived from clicks, and as a result, it may fail to secure any clicks or sales. This bias undermines the reliability of negative relevance signals that are based solely on clicks or sales.
@@
The interaction between the Sellers, Advertising and Search displays an inherent imbalance — Sellers have the ultimate say in adopting Advertising’s keyphrases while Search has the ultimate say in rejecting those keyphrases at the time of auction. For instance, take Itm1 from Figure 3; Advertising considers KP1 and KP4 irrelevant, while eBay Search regards KP2, KP3, and KP5 as irrelevant and the seller rejecting KP2 as a legit keyphrase, leading to Itm1 not participating in any auctions. This outcome occurs regardless of the true relevance of the keyphrases, which means that even if KP2, KP3, and KP5 are genuinely relevant, Itm1 will not enter those auctions because the seller and eBay Search do not find them relevant. An ideal scenario of complete agreement would be for Itm5 and Itm3 where all parties are in complete agreement.  [^1] For Itm3 the outcome is not as great as Itm5 as there are still some keyphrases which are not getting into auctions. However, that is more of a retrieval problem than a relevance problem, i.e., why are the keyphrases KP2and KP4 getting recommended in the first place?
Understanding the dynamics and mechanisms of auctions is essential for the successful execution of advertising campaigns. However, this understanding is further complicated by the complex nature of sellers’ perceptions. Promoted Listings Priority, our product, enables manual targeting, which allows sellers to select their preferred keyphrases while also granting them the option to follow our recommendations. For instance, should sellers judge KP2 for Itm1 as irrelevant, they may choose to disregard this particular keyphrase, thereby rendering our suggestions ineffective. This scenario can escalate, as sellers often encounter an abundance of "irrelevant" keyphrases, discouraging them from proceeding with the adoption of our product. Such outcomes not only lead to diminished Seller Satisfaction scores, but also represent a wasted utilization of resources and opportunities for eBay. Moreover, offering ostensibly "sensible" keyphrases lacking substantial targeting capabilities can result in subpar advertiser campaign performance, potentially prompting sellers to leave the platform. Consequently, it is imperative to ensure that keyphrases resonate with human judgment while simultaneously maintaining robust performance. In this study, we interpret Advertiser keyphrase relevance as a multifaceted interplay among three dynamical systems: seller judgment impacting their product adoption, Advertising supplying the bid and the keyphrases, and Search orchestrating auctions for those keyphrases.
@@ 3.1.2 Click Data
Since clicks are a strong indicator for relevance, we gather click data amounting to roughly 24 million item keyphrase pairs over a period of 30 days with at least 30 impressions and 1 clicks with a minimum CTR of 0.1. Click data can be seen as data annotated by millions of buyers, and it is a stronger signal as buyers are spending their time to do so. The minimum thresholds of CTR, impressions and clicks are used to eliminate noise such as 1 click 1 impression or 1 click and 1000 impression. An important thing to note is that due to the nature of the auction process, all click data is deemed relevant by eBay Search (i.e., they are only surfaced to buyers because it is deemed relevant by eBay Search).